# Bounded domains on Kobayashi hyperbolic manifolds covering compact complex manifolds

### v1

@@ 1 Introduction
Let $M^{n}$ be a (Kobayashi) hyperbolic complex manifold, and let $\Omega\subset M$ be a subdomain with nonempty boundary, and assume $\partial\Omega$ is $C^{1,1}$. Suppose there exists a totally real boundary point $p\in\partial\Omega$. Suppose further that $\Omega$ covers a compact complex manifold. Then $\Omega$ is biholomorphic to the Euclidean ball in $\mathbb{C}^{n}$.
@@
Our proof follows much of the same logic as Zimmer’s. The difference in the arguments is in the construction of the rescaling maps. For this construction, we use a localization technique of Lin and Wong [12] to force the domain $\Omega\subset M$ to be considered as a domain in $\mathbb{C}^{n}$. This construction alters slightly the statements of Zimmer’s results. For this reason, we note that our argument is, in some places, identical to Zimmer’s, which we indicate in the presentation of our proof.
@@ 2.1 Hyperbolic Manifolds
For a domain $\Omega\subset\mathbb{C}^{n}$, the Kobayashi-Royden pseudometric is defined by
$k_{\Omega}(x,v)=\inf\left\{\frac{1}{\lambda}\mid f\in\textup{Hol}(\mathbb{D},\Omega),f(0)=x,f^{\prime}(0)=\lambda v,\lambda>0\right\}.$
We let $d_{\Omega}^{K}(x,y)$ denote the induced Kobayashi pseudodistance of $\Omega$. It has the explicit form
[…]
Here, $z,w\in\Omega$ and $\gamma:[0,1]\to\Omega$ is a curve starting at $z$ and ending at $w$.
@@
As it is pertinent to our main result, we mention a result concerning the Kobayashi metric on bounded domains. Recall that a proper metric space is one in which bounded sets are relatively compact.
@@ 2.2 Rescaling
The Kobayashi metric is continuous with respect to the local Hausdorff topology.
@@
Suppose $\Omega_{j}\subset\mathbb{C}^{n}$ is a sequence of convex domains and $\Omega=\lim_{j\to\infty}\Omega_{j}$ in the local Hausdorff topology. Assume the Kobayashi metric is nondegenerate on $\Omega$ and each $\Omega_{j}$. Then
@@
Fix $r>0$, a sequence $r_{j}>0$ converging to 00, and the sequence of linear maps
@@ 2.3 Bergman Kernel
The computations necessary for the proof of Theorem 1.1 require the following well-known description of the Bergman kernel. See [11] for more details and examples.
@@
[…]
@@
The convergence of the series is absolute and uniform on compact subsets of $\Omega\times\Omega$, and for any $z\in\Omega$, the diagonal $\kappa_{\Omega}(z,z)$ is strictly positive.
The next result - Proposition 2.9 - is typically referred to as the monotonicity property. A proof can be found in any standard several complex variables text. See, for instance, [11].
@@ 2.4 Bounded Symmetric Domains
Definition 2.13.
Let $\Omega\subset\mathbb{C}^{n}$ be a bounded symmetric domain.
  1. (1)
    There is an embedding $F:\Omega\hookrightarrow\mathbb{C}^{n}$ whose image is convex and bounded, called the Harish-Chandra embedding. We denote the image of the embedding by $F(\Omega)=\Omega_{HC}$.
  2. (2)
    We say $\Omega$ is in standard form if it coincides with the image of its Harish-Chandra embedding, i.e., $\Omega=\Omega_{HC}$.
@@
Let $X$ be a compact complex manifold with $c_{1}(X)<0$, and let $\widetilde{X}$ be the universal cover. If $\textup{Aut}\left(\widetilde{X}\right)$ is nondiscrete, then $\widetilde{X}$ is biholomorphic to either
@@ 3 A Localization Technique
Let $\Omega$ be a domain on a complex manifold $M$, and let $p\in\partial\Omega$ be a fixed boundary point. A boundary neighborhood of $p$ is an open set $\widehat{\Omega}=\Omega\cap U$, where $U$ is an open set in $M$ containing $p$.
@@
We note here that, by construction, a boundary neighborhood is biholomorphic to a bounded domain in $\mathbb{C}^{n}$.
@@
Lemma 3.7.
Let $\Omega$ be a domain on a hyperbolic complex manifold admitting a compact quotient. Then there exists a compact set $K\subset\Omega$ such that for every $y\in\Omega$ there is $t\in K$ and $g\in\textup{Aut}(\Omega)$ such that $g(t)=y$ (i.e., $\textup{Aut}(\Omega)\cdot K=\Omega$).
@@
Lemma 3.9.
Let $\Omega$ be a domain on a hyperbolic complex manifold $M$, and suppose there is a totally real boundary point $p\in\partial\Omega$. Let $\{m_{j}\}\subset\textup{Aut}(\Omega)$ be a sequence such that $m_{j}(x)\to p$ for some $x\in\Omega$. Then for any compact set $K\subset\Omega$ and any boundary neighborhood $\widehat{\Omega}$ of $p$, $m_{j}(K)\subset\widehat{\Omega}$ for sufficiently large $j$. In particular, $m_{j}(y)\to p$ for any $y\in\Omega$.
@@ 4.1 Rescaling the Domain
Lemma 4.1.
Let $\Omega\subset\mathbb{C}^{n}$ be a bounded domain with $C^{1,1}$ boundary. After applying an affine transformation, we can assume that
@@
for every $p\in\partial\Omega$. So, since $\Omega\subset B_{1}(e_{1})$, we have that $n_{\Omega}(0)=e_{1}$. Thus $B_{r}(re_{1})\subset\Omega$. ∎
@@
Let $M^{n}$ be a (Kobayashi) hyperbolic complex manifold, and let $\Omega\subset M$ be a subdomain with nonempty boundary, and assume $\partial\Omega$ is $C^{1,1}$. Suppose there exists a totally real boundary point $p\in\partial\Omega$. Suppose further that $\Omega$ covers a compact complex manifold. Then $\Omega$ is biholomorphic to a domain $D\subset\mathbb{C}^{n}$ such that
@@
Suppose that $M^{n}$ is a hyperbolic complex manifold. Let $\Omega\subset M$ be a bounded domain with $C^{1,1}$ boundary admitting a compact quotient. Let $K\subset\Omega$ be the fundamental set of $\textup{Aut}(\Omega)$, i.e., $K$ is compact and $\textup{Aut}(\Omega)\cdot K=\Omega$. Let $p\in\partial\Omega$ be totally real, and let $\widehat{\Omega}$ be a boundary neighborhood of $p$. Note that $\widehat{\Omega}$ is biholomorphic to a bounded domain $\widetilde{\Omega}\subset\mathbb{C}^{n}$. Let $\phi$ be the biholomorphism.
@@
for some $r\in(0,1)$. By Lemma 3.9, there exists a sequence $\{m_{j}\}\subset\textup{Aut}(\Omega)$ such that for any compact set $L\subset\Omega$, $m_{j}(L)\subset\widehat{\Omega}$ for sufficiently large $j$. Fix a sequence $p_{j}\in\widehat{\Omega}$ converging to $p$, and a sequence $k_{j}\in K$ such that $m_{j}(k_{j})=p_{j}$.
Next, fix a sequence $r_{j}\in(0,r)$ converging to 00, and pick $g_{j}\in\textup{Aut}(\widetilde{\Omega})$ such that $g_{j}(\phi(p_{j}))=r_{j}e_{1}$. Consider the $\widetilde{\Omega}$-dilations
@@
Let $\widetilde{\Omega}_{j}:=\Lambda_{j}(\widetilde{\Omega})$ and $F_{j}:=\Lambda_{j}\circ g_{j}\circ\phi\circ m_{j}:\Omega\to\widetilde{\Omega}_{j}$. Then
@@ 4.2 The One-Parameter Subgroup
$\liminf_{j\to\infty}d_{\widetilde{\Omega}_{j}}^{K}(e_{1},z_{j})<\infty,$
@@
$B_{\varepsilon}(w_{0})\subset F_{j}(\Omega)=\Lambda_{j}(\widetilde{\Omega})$
@@
$W:=\left\{w_{0}+(x+iy)e_{1}\ \Bigg{|}\ -\frac{\varepsilon}{2}\leq x\leq\frac{\varepsilon}{2},-m\varepsilon\leq y\leq m\varepsilon\right\}\subset\Lambda_{j}(\widetilde{\Omega}).$
@@ 4.3 Biholomorphic to the Ball
Let $\Omega$ be a subdomain of a hyperbolic complex manifold. Suppose $\partial\Omega$ is $C^{1,1}$, and that there exists a totally real boundary point $p\in\partial\Omega$. If $\Omega$ covers a compact manifold, then $\Omega$ is biholomorphic to a bounded symmetric domain in $\mathbb{C}^{n}$.
@@
The proof of Proposition 4.9 again exploits the geometry of the rescaled domain $D$, and makes use of the theory of bounded symmetric domains. Zimmer introduces a holomorphic function which measures the volume contraction (or expansion) of the biholomorphism from Proposition 4.3 along a linear slice of the rescaled domain, $D$. A contradiction argument on the real rank of $\Omega$ as a bounded symmetric domain completes the proof. A key part of the argument is showing that one can parameterize the diagonal of a maximal polydisk in the Harish-Chandra embedding of $\Omega$ via the Jacobian matrix of the biholomorphism. Propositions 2.9 and 2.10 are instrumental here.
@@ 5 Conclusions
    $M=\widetilde{M}/\Gamma$, where $\widetilde{M}$ is the universal cover of $M$ and $\Gamma$ is a discrete subgroup of the identity component of $\textup{Aut}\left(\widetilde{M}\right)$ acting freely on $\widetilde{M}$.
@@
Proposition 1.2 and, or course, Zimmer’s paper [21] were the inspiration for our method of proof.
@@ Acknowledgement
The author thanks Professor Bun Wong for his invaluable guidance during the early stages of this research.

### v2

@@ 1 Introduction
Let $M^{n}$ be a taut complex manifold, and let $\Omega\subset M$ be a subdomain with nonempty boundary, and assume $\partial\Omega$ is $C^{1,1}$. Suppose there exists a totally real boundary point $p\in\partial\Omega$. Suppose further that $\Omega$ covers a compact complex manifold. Then $\Omega$ is biholomorphic to the Euclidean ball in $\mathbb{C}^{n}$.
@@
Our proof follows much of the same logic as Zimmer’s. The main difference in the arguments is in the construction of the rescaling maps. For this construction, we use a localization technique of Lin and Wong [12] (see Section 3) to force the domain $\Omega\subset M$ to be considered as a domain in $\mathbb{C}^{n}$. In their paper, Lin and Wong accomplish this by choosing holomorphic coordinates near a boundary point so that the local geometry is Euclidean. The boundary then assumes a normalized form. After an additional scaling, the localized domain converges to a domain in $\mathbb{C}^{n}$. Normal family arguments and other classical tools from several complex variables are now applicable to the original manifold. This construction is crucial in our proof of Theorem 1.1, particularly in the proof of Proposition 4.3. In this result, we introduce new sequences of automorphisms in order to use the localization technique, and so our argument requires careful verification that Zimmer’s computations remain valid.
Since we consider domains of a general complex manifold as opposed to domains in $\mathbb{C}^{n}$, it is natural to expect some differences between the hypotheses of our main result and Zimmer’s (Theorem 1.3). In particular, the tautness assumption is necessary in order to use the localization technique. Some of these differences are slight, and so we note that our argument is, in some places, identical to Zimmer’s, which we indicate in the presentation of our proof.
@@ 2.1 Hyperbolic Manifolds
Let $M$ be a complex manifold. For $x\in M$ and $v\in T^{(1,0)}M$, the Kobayashi-Royden pseudometric is defined by
$k_{M}(x,v)=\inf\left\{\frac{1}{\lambda}\mid f\in\textup{Hol}(\mathbb{D},M),f(0)=x,f^{\prime}(0)=\lambda v,\lambda>0\right\}.$
We let $d_{M}^{K}(x,y)$ denote the induced Kobayashi pseudodistance of $M$. It has the explicit form
$d_{M}^{K}(z,w)=\inf_{\gamma}\int_{0}^{1}k_{M}(\gamma(t),\gamma^{\prime}(t))dt.$
Here, $z,w\in M$ and $\gamma:[0,1]\to M$ is a piecewise $C^{1}$ curve starting at $z$ and ending at $w$.
@@
As it is pertinent to our main result, we mention a result concerning the Kobayashi distance on bounded domains. Recall that a proper metric space is one in which bounded sets are relatively compact.
@@
Definition 2.4.
A complex manifold $M$ is said to be taut if for every complex manifold $N$, the set of all holomorphic functions from $N$ to $M$ is a normal family.
Remark 2.5.
It follows from the definitions that taut manifolds are always hyperbolic.
@@ 2.2 Rescaling
The Kobayashi distance is continuous with respect to the local Hausdorff topology.
@@
Suppose $\Omega_{j}\subset\mathbb{C}^{n}$ is a sequence of convex domains and $\Omega=\lim_{j\to\infty}\Omega_{j}$ in the local Hausdorff topology. Assume the Kobayashi distance is nondegenerate on $\Omega$ and each $\Omega_{j}$. Then
@@
Fix $r>0$, a sequence $r_{j}>0$ converging to 0, and the sequence of linear maps
@@ 2.3 Bergman Kernel
The computations necessary for the proof of Theorem 1.1 require the following well-known description of the Bergman kernel. See Section 1.4 in [11] for more details and examples.
@@
$H^{2}(\Omega)=\left\{f\in\textup{Hol}(\Omega,\mathbb{C})\ \bigg|\ \int_{\Omega}|f|^{2}d\mu<\infty\right\}.$
@@
The convergence of the series is absolute and uniform on compact subsets of $\Omega\times\Omega$, and for any $z\in\Omega$, the diagonal $\kappa_{\Omega}(z,z)$ is strictly positive. Additionally, $\kappa_{\Omega}$ does not depend on the choice of orthonormal basis.
The next result - Proposition 2.11 - is typically referred to as the monotonicity property.
@@ 2.4 Bounded Symmetric Domains
Let $\Omega\subset\mathbb{C}^{n}$ be a bounded symmetric domain. Harish-Chandra constructed an embedding $F:\Omega\hookrightarrow\mathbb{C}^{n}$ whose image is convex and bounded (see [16], Chapter II, Section 4). Moreover, there exists a norm $\|\cdot\|_{HC}$ on $\mathbb{C}^{n}$ such that
$F(\Omega)=\left\{z\in\mathbb{C}^{n}\ \mid\ \|z\|_{HC}<1\right\}.$
We denote the image of this embedding by $F(\Omega)=\Omega_{HC}$. We say that $\Omega$ is in standard form if it coincides with the image of its Harish-Chandra embedding, i.e., $\Omega=\Omega_{HC}$.
@@
[…]
@@ 3 A Localization Technique
Let $\Omega$ be a domain on a complex manifold $M$, and let $p\in\partial\Omega$ be a fixed boundary point. A boundary neighborhood of $p$ is an open set $\widehat{\Omega}=\Omega\cap U$, where $U$ is a relatively compact coordinate chart centered at $p$.
@@
Since $U$ is a relatively compact coordinate chart, its image under the chart map is a bounded open subset of $\mathbb{C}^{n}$. Therefore, a boundary neighborhood is biholomorphic to a bounded domain in $\mathbb{C}^{n}$.
@@
Lemma 3.7 (Lemma 2.1 in [12]).
Let $\Omega$ be a domain on a taut manifold admitting a compact quotient. Then there exists a compact set $K\subset\Omega$ such that for every $y\in\Omega$ there is $t\in K$ and $g\in\textup{Aut}(\Omega)$ such that $g(t)=y$ (i.e., $\textup{Aut}(\Omega)\cdot K=\Omega$).
@@
Lemma 3.9 (Lemma 2.4 in [12]).
Let $\Omega$ be a domain on a taut manifold $M$ admitting a compact quotient. Let $\{x_{j}\}\subset\Omega$ be a sequence converging to a boundary point $p\in\partial\Omega$. Then there exists a sequence $\{m_{j}\}\subset\textup{Aut}(\Omega)$ such that $m_{j}^{-1}(x_{j})$, through a subsequence if necessary, converges to a point $x\in\Omega$. Moreover, $m_{j}(x)\to p$.
Lemma 3.10 (Lemma 2.5 in [12]).
Let $\Omega$ be a domain on a taut manifold $M$, and suppose there is a totally real boundary point $p\in\partial\Omega$. Let $\{m_{j}\}\subset\textup{Aut}(\Omega)$ be a sequence such that $m_{j}(x)\to p$ for some $x\in\Omega$. Then for any compact set $K\subset\Omega$ and any boundary neighborhood $\widehat{\Omega}$ of $p$, $m_{j}(K)\subset\widehat{\Omega}$ for sufficiently large $j$. In particular, $m_{j}(y)\to p$ for any $y\in\Omega$.
@@ 4.1 Rescaling the Domain
Lemma 4.1 (Lemma 4.5 in [22]).
Let $\Omega\subset\mathbb{C}^{n}$ be a bounded domain with $C^{1,1}$ boundary. After applying an affine transformation, we can assume that $0\in\partial\Omega$ and
@@
for every $p\in\partial\Omega$. So, since $0\in\partial\Omega$, by assumption, we have that $n_{\Omega}(0)=e_{1}$. Thus $B_{r}(re_{1})\subset\Omega$. ∎
@@
Let $M^{n}$ be a taut manifold, and let $\Omega\subset M$ be a subdomain with nonempty boundary, and assume $\partial\Omega$ is $C^{1,1}$. Suppose there exists a totally real boundary point $p\in\partial\Omega$. Suppose further that $\Omega$ covers a compact complex manifold. Then $\Omega$ is biholomorphic to a domain $D\subset\mathbb{C}^{n}$ such that
@@
Suppose that $M^{n}$ is a taut manifold. Let $\Omega\subset M$ be a bounded domain with $C^{1,1}$ boundary admitting a compact quotient. Let $K\subset\Omega$ be the fundamental set of $\textup{Aut}(\Omega)$, i.e., $K$ is compact and $\textup{Aut}(\Omega)\cdot K=\Omega$. Let $p\in\partial\Omega$ be totally real, and let $\widehat{\Omega}$ be a boundary neighborhood of $p$. Note that $\widehat{\Omega}$ is biholomorphic to a bounded domain $\widetilde{\Omega}\subset\mathbb{C}^{n}$ (see Remark 3.2). Let $\phi$ be the biholomorphism.
@@
for some $r\in(0,1)$. Fix a sequence $p_{j}\in\widehat{\Omega}$ converging to $p$. By Lemma 3.9, there exists a sequence $\{m_{j}\}\subset\textup{Aut}(\Omega)$ such that $m_{j}(x)\to p$ for some $x\in\Omega$. Then by Lemma 3.10, for any compact set $L\subset\Omega$, $m_{j}(L)\subset\widehat{\Omega}$ for sufficiently large $j$. In particular, this holds for the fundamental set $K$. So, fix a sequence $k_{j}\in K$ such that $m_{j}(k_{j})=p_{j}$.
Next, fix a sequence $r_{j}\in(0,r)$ converging to 0. Since $\textup{Aut}(B_{1}(e_{1}))$ acts transitively on $B_{1}(e_{1})$, we may pick $g_{j}\in\textup{Aut}(B_{1}(e_{1}))$ such that $g_{j}(\phi(p_{j}))=r_{j}e_{1}$. The restriction of $g_{j}$ to $\widetilde{\Omega}$ is a biholomorphism from $\widetilde{\Omega}$ onto the domain $g_{j}(\widetilde{\Omega})\subset B_{1}(e_{1})$. Consider the $\widetilde{\Omega}-$dilations
@@
[…]
@@ 4.2 The One-Parameter Subgroup
$\liminf_{j\to\infty}d_{\widetilde{\Omega}_{j}}^{K}(e_{1},z_{j})<\infty.$
@@
[…]
@@
[…]
@@ 4.3 Biholomorphic to the Ball
Let $\Omega$ be a subdomain of a taut manifold. Suppose $\partial\Omega$ is $C^{1,1}$, and that there exists a totally real boundary point $p\in\partial\Omega$. If $\Omega$ covers a compact manifold, then $\Omega$ is biholomorphic to a bounded symmetric domain in $\mathbb{C}^{n}$.
@@
The proof of Proposition 4.9 again exploits the geometry of the rescaled domain $D$, and makes use of the theory of bounded symmetric domains. Zimmer introduces a holomorphic function which measures the volume contraction (or expansion) of the biholomorphism from Proposition 4.3 along a linear slice of the rescaled domain, $D$. A contradiction argument on the real rank of $\Omega$ as a bounded symmetric domain completes the proof. A key part of the argument is showing that one can parameterize the diagonal of a maximal polydisk in the Harish-Chandra embedding of $\Omega$ via the Jacobian matrix of the biholomorphism. Propositions 2.11 and 2.12 are instrumental here.
@@ 5 Conclusions
[…]
@@
Proposition 1.2 and, of course, Zimmer’s paper [22] were the inspiration for our method of proof.
@@ Acknowledgement
The author thanks Professor Bun Wong for his invaluable guidance during the early stages of this research. The author also extends his gratitude to the anonymous reviewer for their detailed suggestions and observations.